OpenStudy (anonymous):

Anyone good at 8th grade math ?

OpenStudy (anonymous):

@Daniee_Bruhh

OpenStudy (anonymous):

It depends on the question.

OpenStudy (anonymous):

@Jaynator495 @TheSmartOne

OpenStudy (anonymous):

i'm in 8th grade and i can try to help u

RhondaSommer (rhondasommer):

I got you

OpenStudy (anonymous):

what's the question

OpenStudy (anonymous):

The selected one is what i think it is

RhondaSommer (rhondasommer):

correct

OpenStudy (anonymous):

@DIDITHEDAD

OpenStudy (anonymous):

easy

OpenStudy (anonymous):

your right

RhondaSommer (rhondasommer):

:)

OpenStudy (anonymous):

There are 6 more @RhondaSommer @studygirl13

RhondaSommer (rhondasommer):

6x - 2x is what

OpenStudy (anonymous):

12x

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

4x

RhondaSommer (rhondasommer):

yep

RhondaSommer (rhondasommer):

so \[4x-5=-6+21\] 21 - 6 is what

OpenStudy (anonymous):

15

OpenStudy (anonymous):

So its d ????

RhondaSommer (rhondasommer):

\[4x-5=15\]add 5 to both sides of the equation

RhondaSommer (rhondasommer):

yep :)

OpenStudy (anonymous):

Now this one and the next two after it, the selected one is what i think the answer is

RhondaSommer (rhondasommer):

tag me in the next question please

OpenStudy (anonymous):

hey